bill533:

A circle is centered at C(−1,−3) and has a radius of 6.Where does the point P(-6,-6) lie? A. Inside the circle B. on the circle C. outside the circle

sillybilly123:

Check the distance "\(\Delta\)" from C to P: \(\Delta^2 = (-6 -(- 1))^2 + (-6 -(- 3))^2\) \( = 25 + 9 = 34 \) Whereas \(R^2 = 36\)